OpenStudy (alanna):

Hi, I'm confused how to solve: 3x^2-2x-(x^2-5x), thanks!

OpenStudy (anonymous):

=3x^2-2x-x^2+5x =2x^2+3x btw, why? this is easy

OpenStudy (anonymous):

=3x^2-2x-x^2+5x =2x^2+3x btw, why? this is easy

OpenStudy (alanna):

I already have the answer, its a practice quiz, but unfortunately that doesn't help me understand how to do it even though its easy

OpenStudy (anonymous):

you have a minus sign outside the bracket, that means (-1) times everything in bracket, so -(x^2-5x) = -1 * (-x^2-5x) = (-1)*(x^2)+(-1)*(-5x)=-x^2+5x so ur equation becomes 3x^2-2x-x^2+5x=2x^2+3x hope that help?

OpenStudy (anonymous):

sorry, no minus before x^2 in the bracket in the second one from the left :)

OpenStudy (alanna):

Thanks!

OpenStudy (alanna):

Thanks!